# The Physics and Metaphysics of Social Powers: Bridging Cognitive Processing and Social Dynamics, a New Perspective on Power through Active Inference

### v1

@@ Abstract
Empowered individuals attract the attention of others, who contribute to increasing the scale of their access to various policies effectuating these state spaces. The presented argument posits that social power, in the context of active inference, is a function of several variables, among them: an individual’s or group’s ability to attract the attention of others with the ability to facilitate desired policies, and an individual’s or group’s capacity to effectively process information. As a result of its power-amplifying effects, this extended computational ability also buffers against possible vulnerabilities. Semantic social artifacts (narratives, ideologies, histories, representation, etc.) as attentional scripts (ensuring the coordination of social patterns: psychological schemata, habits, groupthink, insulation, policies, etc.: Albarracin et al., 2021), are the structures under and by which agents mediate and leverage power, and are thus this article’s main objects of analysis.
@@
Narratives are social mechanisms which have the ability to make or break power (Blakeley, 2024). Not-yet-engaged scripts (speculations, fictions, etc.: future forecasting) can be understood as semantic attractors enabling—relatively—risk-free learning (de Jager, 2022). Possibilistic cognitive explorations in the form of, e.g., (fictional) narratives allow agents to consider counterfactual scenarios and minimize uncertainty without exposure to “real-world” consequences. This mechanism frames both the evolutionary advantage and persistent appeal of narrative engagement in human culture. In the context of power, leveraging “strong narratives”—such as “us versus them” reductionisms—greatly minimizes uncertainty (technically: variational free energy, the divergence between what is expected and what is actually experienced) about an agent’s or group’s current predicaments, and thus become appealing (sometimes self-fulfilling) policies that bring about their effects, resulting in collective behaviors which can be beneficial or detrimental to the greater whole (Albarracin et al., 2022).
@@
The effects of variational free energy-minimization can be observed at the level of the basic structures of language: (semantic) attractors such as redundancy, rhyme and alliteration, can be understood as patterns which aid memory, communication and information-retrieval policies: they permanently reappear, pooling attention to the “same” phenomena, thus greatly reducing the need to decode and recode linguistic interaction anew. By way of the same uncertainty-minimizing mechanisms, (simple or complex) narrative scripts create cognitive situations where agents can reduce complex aspects of the world down to preferred and/or already-identified scenarios. Unlike “direct” experience, where prediction errors present immediate risks, narratives provide speculative semantic spaces for exploring models against hypothetical outcomes. With regard to social power, the problems immediately emerge in situations where collective attention controlled by agents already benefiting from dominant narratives can lead to, e.g., the exacerbation of misinformation, echo chambers, conspiracy theories, etc. (ibid.), in an attempt to retain attention and therefore power. In political communication and organization, where the proliferation of scripts can be understood as an optimization of the learning/risk trade-off inherent in active inference, this is of crucial attention. Particularly at a time when the democratic, technological and futurological organization of life is in question.
@@
“… power must be understood in the first instance as the multiplicity of force relations immanent in the sphere in which they operate and which constitute their own organization; as the process which, through ceaseless struggles and confrontations, transforms, strengthens, or reverses them; as the support which these force relations find in one another, thus forming a chain or a system, or on the contrary, the disjunctions and contradictions which isolate them from one another; and lastly, as the strategies in which they take effect…” (1990 [1978], pp. 92–93, our emphasis in bold).
@@
The phenomenon, as well as the representations of what precisely constitutes the social acquisition and maintenance of power, is thus a complex series of multidirectional processes which cannot be simplistically or hierarchically schematized, but must be analyzed in terms of interactions between agents and the constraint-regimes they are embedded within. Traditionally, the framing of socially-facilitated empowerment—such as that of the state, or of a large corporation—has been understood as the effectuation of (often top-down) strategies, tactics, policies, narratives (i.e., scripts, as will be argued later) which grant empowered agents/systems more leverage, that is: more influence on future states 111 Note: here we refer to the meaning of state as “condition”, not in the political sense. However, the double entendre is opportune in this context. . We will frame this in terms of information-processing capacities and opportunities, where powerful agents either rely on others to bring about desired policies (i.e., exploitation: in the biological and sociological senses), or enjoy more information-processing possibilities as a result of relying on others for the reproduction of (material) tasks (i.e., exploration). The effects of social empowerment therefore imply an increased ability to compute information, thereby augmenting the evolution of a specific state space. Basing our arguments in active inference, as will be explained, will allow for an analysis of “[p]ower [a]s the probability that one actor within a social relationship will be in a position to carry out [their] will despite resistance,” (Weber, (1925) 1965, p. 152) in terms of actual  [^2] Weber’s quote ends in “… regardless of the basis on which this probability rests.” (ibid.). We will challenge this through a material-informatic analysis, as the basis of this probability is not trivial. See, e.g., Ramstead et al., 2023, on the introduction of a “formal language for modelling the constraints, forces, potentials, and other quantities determining the dynamics of such systems, especially as they entail dynamics on a space of beliefs (i.e., on a statistical manifold)” for more background on the possible bridges between material and informatic in terms of Bayesian mechanics. Also note that the translation of “probability” in Weber has also been that of “ability”, “capacity”, and other terms, we take one of the most popular formulations here which also aids our purpose of framing this in terms of a probabilistic mechanics. probabilities. We will present novel understandings of concepts such as scripts and attention, by analyzing power in terms of the physics of beliefs (Ramstead et al., 2023), and the metaphysics of social organization, where metaphysics is understood as the possibilistic exploration of state spaces (i.e., the future).
@@
The outline of the article is as follows: in the subsections that follow we present a necessarily broad and brief elucidation of different understandings of power, primarily by way of philosophy and sociology, in order to set the focus on the concept of power as the control of attention via explorative and exploitative scripts, which can exist as shared metacognitive computing conditions (i.e., agents are relatively aware of the implications of scripts), coercive domination (where agents are subsumed under a narrative which does not benefit them), or: often an inevitable combination of both. In section 2 we explain the formal context of Active Inference as the theoretical ground which will sustain our analysis. In section 3 we synthesize the previous sections towards an explanation of social scripts and power, and how scripts can frame the concept of narratives, attention, coercion and resistance. In section 4 we look at power through the lens of Active Inference, and in section 5 we treat the formal framework integration. We conclude by presenting attentional strategies which can result in alternative modulations of power.
@@ 0.1 Power-pooling in, through and between agents
Gravitational metaphors abound in language (Shepard 1987, Lakoff and Johnson (1980) 2003, Dennett 1992, de Jager 2021, Kent 2024333 Lachlan Kent (2024) has further argued that the vestibular system, which regulates graviception, also affects the higher-order consciousness of selfhood: gravity constrains physical proprioception and action, as well as the very sense of a narrative self (Dennett 1992). ). Such gravitational metaphors find analogues in the network sciences, whereby preferential attachment processes lead to additional power accruing increasingly to those already in possession of it (Barabási and Albert 1999). The idea of power being pooled towards complex social attractors was famously presented by Pierre Bordieu, who linked various concepts of capital (cultural, political, etc.) to power as social (at)traction, borrowing the metaphor from physical field theory [^4] Just as electromagnetic fields exert forces on particles depending on their position and properties, Bourdieu conceptualized social fields as spaces where agents’ positions are (pre)determined by their access to different forms of capital (economic, cultural, symbolic, etc.). Bourdieu’s use of field theory therefore frames system structure and the dynamics of agency: a given field structures possible actions by defining what counts as accepted or legitimate, and habitus as inclinations; dispositions, determine how fields are navigated: possibilities. The ensuing social practices herein are the result of heuristics comparable to satisficing/fast and frugal heuristics, which in our context are formalizable through AIF.. We expand on his legacy which takes the ontoepistemological and the empirical to be inseparable—and his desire to ground sociological analyses of habitus in a scientifically tractable manner—by framing these attentional attractors that move systems toward specific configurations, through Active Inference (AIF). Another key thinker of power in the 20th century was Hannah Arendt. Arendt’s relational idea of collective power frames its essence as communicative, public and permanently potential: people may gain momentum together, but power is never guaranteed. Coercion, (symbolic) violence and oppressive authority, in her framing, do not constitute power but are rather reductions of its possibility. In our context, this differentiation would imply the first presentation being shared cultural narratives towards explorative change, and the second as top-down exploitative practices which reduce the possibility of novelty in given structures (systems such as communities; identities, histories, etc.). Similarly, more recently, Grace Blakeley has analyzed the disempowerment of collective political imagination by presenting this as the disillusion555 In our terms, perhaps, understood as the loss of attentional momentum. with shared narratives: democratic engagement in the context of, e.g., British politics, has been affected by a sense of lack of agency (from “there is no alternative” all the way to the not-so “free” market666“This sense of unfreedom is grounded in the deep disparities of power that exist within capitalist societies—many of which are completely invisible. Most people are denied autonomy over their lives, yet we are consistently told that we are free to choose how we live. Life under capitalism means life under a system in which decisions about how we work, how we live, and what we buy have already been made by someone else. Life under capitalism means living in a planned economy, while being told that you are free.” (Blakeley 2024, p. 6, our emphasis in bold).). This resulting political disenchantment was framed earlier by Mark Fisher in terms of the privatization or individualization of systemic stress (Fisher, 2009): through strategic narratives which set public attention on individual responsibility, publics become overwhelmed by having to resolve, at the individual level, stress which is in fact systemic. Whether we are talking about a metastabilizing “invisible hand” or modeling swarm-organization (Wareham et al., 2023): we know of many systems that have infelicitous critical points when scaling: requisite variety demands that control systems maintain sufficient internal complexity (entropy) to match the complexity of what they aim to control. This can be understood at a number of levels (in modeling: computational tractability problems regarding communication between agents that have to do with temporal tracking and issues such as latency, for example), but at the level of social dynamics and the (in)efficiencies of power asymmetries, this is most obvious in the context of the power of attention: the assertion of a specific narrative, if it gains traction as a collective belief, will result in the simplification (and mobilization) of desire. That is: a strong-enough narrative can become a script which agents incorporate as part of their generative model, rather than as a meta-consideration (a highly possibilistic variable), existing on a dialogical landscape. In crude terms: the difference between the first and the second can be exemplified by the situation where an agent believes that any single solution can solve national scale problems, such as the deportation of immigrants supposedly solving every socio-economic issue in a very large nation. Our framing allows for a new take on the dialectics between the individual and the collective. In terms of questions of information-processing capacity: not only are resource-distribution and the division of labor a systemic fact, but also: since the modulation and projection of future possibilities cannot be computed at the level of the individual due to its high complexity, this is how and why the minimization of uncertainty for social systems is, in fact, social. Agents seek alliances and policies which ensure cognition (i.e., information processing and future-projection) can be distributed (Vasil et al., 2020), as this allows for efficient practical traction because of “epistemic confidence in knowing, interpreting, and acting together” (Shteynberg et al., 2023). This is the basis for any division of labor, where specialized roles and collaborative synergy optimize task execution. As demonstrated by Ispolatov et al. (2011), division of labor in biological systems improves fitness by compartmentalizing incompatible processes, while Labella (2007) shows its application in robotics, enhancing adaptability and resource efficiency. Marzband et al. (2017) show that coalition-based strategies in distributed energy resources maximize collective benefits, offering a model for scalable and cooperative systems in human and artificial contexts. This is in contrast with individualized agency—which is a prerequisite for social coercion—where one’s estimation of “divergence in mental states between self and other(s)” allows for the “prediction and interpretation of others’ behavior” (Shteynberg et al., 2023), resulting in shorter-range practical effects. Remaining politically-agnostic in any analysis presents us with some questions: what can we assume to actually know about each other when we act collectively? The concept of a ‘self’ as a single agent seems particularly questionable, especially as all selves exist within a regime of constraints which exert effects on their supposed autonomy. A famous response to the Foucauldian take on power and the individual came from Gilles Deleuze. His concept of the “-dividual” replaces the image of, e.g., the factory worker as a subject of disciplinary capitalism, with the image of a distributed subjectivity, e.g., the owner of a credit card as a subject of libertarian debt, occupying various spatiotemporal scales and modulated not by clock-time or physical enclosure, but by probabilities and contractual fixation/limitation (Deleuze, 1990). The latter condition emerges, Deleuze shows, as a result of novel paradigms of technocapitalism which present new constraint regimes. Distribution, collectivity, or shared sameness, therefore, do not always equal autonomous collective traction, but can also be the result of policies which pool power towards attractors, where the agents implicated within them may not even be aware of their implication (the examples in capitalism are vast: we have very little idea of the complexity behind the products we buy into, but we are coerced into buying nonetheless [^7] For more on this theme, see “Post-control Script-societies” (de Jager, 2023).). In the interest of this article, the concept tying all these processes towards an analysis of power dynamics is that of attention. We will briefly treat this below (and in later sections) before moving to section 2 for an exposition of the AIF framework.
@@ 0.2 Attention as power: shared narratives and awareness or coercion
Exerting social control over others’ actions begins at the level of attention: whether this be by attracting attention (“follow me!”), or directing attention (“look at that!”). Our interest in formalizing certain aspects of power dynamics is directing attention towards (metacognitive and collective) scripts with the capacity of empowering responses to counter dominant strategies which, currently, minimize uncertainty for the very few [^8] Again, often by allocating too much importance to the atomized individual, which is a predictably convenient way to simplify the actions of complex distributed processes such as individuals, into overseeable patterns. There is nothing essentially “wrong” with this, this is the way perception tracks reality. However, revealing these dynamics allows us an entrance into the ways in which we may modulate (the effects of) these phenomena/concepts.. Shared structures of meaning can exist as collective metacognitive computing conditions, i.e., conditions in which agents are—relatively—aware of the implications of the scripts they follow, or coercive domination: where agents are subsumed under a script which, often, does not benefit their uncertainty-minimization on the long run. If we understand successful narratives as those which exert the most effect [^9] Often “fast and frugal” (Gigerenzer, Simon) narratives, which are low-risk and, in most political contexts, imply promises or solutions impossible to accomplish: “you have been lied to”, “the other is the problem”, etc., analyzing them in terms of attention seems logical: in the context of uncertainty-minimization, simple narratives have the capacity to dominate vast regimes of attention (economies), as they are scripts which minimize the complexities of life into low-computation, predictable visions [^10] Whether these are dialogical and representational/symbolic or surreptitious and encrypted: we’re dealing with different strategies, but similar results..
@@
In the transition between an economy centered around material and spatial capital, towards one of temporal, speculative capital, we can straightforwardly connect attention to computation by understanding the implicit link between risk-assessment and investment: (possible) future gains ought to be calculated, and directing investment begins with attentional gains111111“A sufficiently large financial institution has the power to direct investment into certain technologies and therefore to determine which futures are available and which ones are foreclosed. And all these firms can consolidate their power by buying up competitors, erecting barriers to entry, and crushing workers’ attempts to organize, further insulating them from competitive pressure and giving them significant authority both within their domains and over society as a whole.” (Blakeley, 2024).. Beyond the strictly financial realm, in the context of futurology and planetary survival, the concept of AI technologies as “powerful” grounds the arguments surrounding possible existential risks. E.g., Cappelen et al., (2024) rely on the concept of “power” in order to assess to what extent humans may or may not lose control over their own technological creations. Likewise, other authors make claims for “powerful yet narrow AI systems” as incapable of accommodating the semantic complexity which grounds mortal life (Ororbia and Friston 2023) [^12] They continue: “One might view the future as residing in not viewing intelligence as a difference engine but instead as an artificial form of sentience: one that is capable of self-healing and self-repairing with autonomy — its existential imperative being to persist in (generalized) synchrony with its world.” (ibid.). The questions of “autonomy” and “synchrony” are precisely the ones we are after in this article, by observing how power, in mortally-determined social systems, is driven by attention.. What these assessments have in common is a concern with attention as a resource: a collectively predicted future is one where attention is directed towards synchronized goals (whether this be investment, alignment, etc.).
@@
Ultimately, because potential power (as attention) is both a material and an informational force, “seizing the means of computation” (Doctorow, 2023) follows in line with a Marxist analysis of capital in our moving paradigm: from materialist analyses of territorial and labor time domination, through a semantic-sociocultural analysis (Bourdieu), to informational analyses where we ought to think about the production of the future by means of projective computation. Power is thus based on the controlling of attention by way of scripts (again: narratives, representation, identities, symbolisms, etc.) and its current substrate in our contemporary landscape is computation: both in the flesh and in silico. As we will see in section 3, power’s dual nature (as force attraction and information computation) in the social realm thus implies the use of scripts as they are the means by which agents gain or lose the ability to direct the attention of other agents (with access to desired policies), and the ability to effectively process information in favor of these policies (and other longer-term possibilities).
@@
The variables modulating regimes of attention that have practical modeling-potential are plenty, and in sections 4 and 5, our interest is in the control, influence, and synchronization of attention, which lead to social power structures. Our focus will be on empowerment quantified as agents’ potential influence over their environment (which is tracked by assessing the mutual information between actions and states). We will analyze this through formal methods that are tractable, but this presents us with indications of future modeling possibilities, which in our context remain unformalized as they are practically intractable. We can think, for example, about the modeling of “inheritance” or “provenance” variables: wherever/however an agent exists, they could be born into a rich family, in a specific part of the world, etc., and these factors have an influence on how others interpret them (and therefore on life-expectancy: mortality, and many other crucial effects). One can also posit a possible “frictions/contingengies/contextualities” variable—in colloquial terms “you have to work with what you’ve got”—by which even if agents can attract others’ attention, and process information effectively towards the effectuation of a specific policy, the projected goal and thus possible state space is constrained by a context which can be highly divergent from the “ideal” one they project as required for bringing a goal about. Confidence in policies is a well-framed and studied variable in the context of AIF, and in the literature it is often directly linked to the concept of power: “self-confidence [is] the person’s experience of their power to act in the world” (Kiverstein et al., 2019). There are more, of course, and it depends on how granular one may want to go, these are examples which give the reader an indication of our intentions in later sections.

### v2

@@ Abstract
Empowered individuals attract the attention of others, who contribute to increasing the scale of their access to various policies effectuating these state spaces. The presented argument posits that social power, in the context of active inference, is a function of several variables, among them: an individual’s or group’s ability to attract the attention of others with the ability to facilitate desired policies, and an individual’s or group’s capacity to effectively process information. As a result of its power-amplifying effects, this extended computational ability also buffers against possible vulnerabilities. Semantic social artifacts (narratives, ideologies, histories, representation, etc.) as attentional scripts (ensuring the coordination of social patterns: psychological schemata, habits, groupthink, insulation, policies, etc.: [Albarracin et al., 2021]), are the structures under and by which agents mediate and leverage power, and are thus this article’s main objects of analysis.
@@
Narratives are social mechanisms which have the ability to make or break power [Blakeley, 2024]. Not-yet-engaged scripts (speculations, fictions, etc.: future forecasting) can be understood as semantic attractors enabling—relatively—risk-free learning [de Jager, 2022]. Possibilistic cognitive explorations in the form of, e.g., (fictional) narratives allow agents to consider counterfactual scenarios and minimize uncertainty without exposure to “real-world” consequences. This mechanism frames both the evolutionary advantage and persistent appeal of narrative engagement in human culture. In the context of power, leveraging “strong narratives”—such as “us versus them” reductionisms—greatly minimizes uncertainty (technically: variational free energy, the divergence between what is expected and what is actually experienced) about an agent’s or group’s current predicaments, and thus become appealing (sometimes self-fulfilling) policies that bring about their effects, resulting in collective behaviors which can be beneficial or detrimental to the greater whole [Albarracin et al., 2022].
@@
The effects of variational free energy-minimization can be observed at the level of the basic structures of language: (semantic) attractors such as redundancy, rhyme and alliteration, can be understood as patterns which aid memory, communication and information-retrieval policies: they permanently reappear, pooling attention to the “same” phenomena, thus greatly reducing the need to decode and recode linguistic interaction anew. By way of the same uncertainty-minimizing mechanisms, (simple or complex) narrative scripts create cognitive situations where agents can reduce complex aspects of the world down to preferred and/or already-identified scenarios. Unlike “direct” experience, where prediction errors present immediate risks, narratives provide speculative semantic spaces for exploring models against hypothetical outcomes. With regard to social power, the problems immediately emerge in situations where collective attention controlled by agents already benefiting from dominant narratives can lead to, e.g., the exacerbation of misinformation, echo chambers, conspiracy theories, etc. (ibid.), in an attempt to retain attention and therefore power. In political communication and organization, where the proliferation of scripts can be understood as an optimization of the learning/risk trade-off inherent in active inference, this is of crucial attention, particularly at a time when the democratic, technological and futurological organization of life is in question.
@@
“… power must be understood in the first instance as the multiplicity of force relations immanent in the sphere in which they operate and which constitute their own organization; as the process which, through ceaseless struggles and confrontations, transforms, strengthens, or reverses them; as the support which these force relations find in one another, thus forming a chain or a system, or on the contrary, the disjunctions and contradictions which isolate them from one another; and lastly, as the strategies in which they take effect…” ([Foucault, 1990], pp. 92–93, our emphasis in bold).
@@
The phenomenon, as well as the representations of what precisely constitutes the social acquisition and maintenance of power, is thus a complex series of multidirectional processes which cannot be simplistically or hierarchically schematized, but must be analyzed in terms of interactions between agents and the constraint-regimes they are embedded within. Traditionally, the framing of socially-facilitated empowerment—such as that of the state, or of a large corporation—has been understood as the effectuation of (often top-down) strategies, tactics, policies, narratives (i.e., scripts, as will be argued later) which grant empowered agents/systems more leverage, that is: more influence on future states 111 Note: here we refer to the meaning of state as “condition”, not in the political sense. However, the double entendre is opportune in this context. . We will frame this in terms of information-processing capacities and opportunities, where powerful agents either rely on others to bring about desired policies (i.e., exploitation: in the biological and sociological senses), or enjoy more information-processing possibilities as a result of relying on others for the reproduction of (material) tasks (i.e., exploration). The effects of social empowerment therefore imply an increased ability to compute information, thereby augmenting the evolution of a specific state space. Basing our arguments in active inference, as will be explained, will allow for an analysis of “[p]ower [a]s the probability that one actor within a social relationship will be in a position to carry out [their] will despite resistance,”(Weber [2009], p. 152) in terms of actual  [^2] Weber’s quote ends in “… regardless of the basis on which this probability rests.” (ibid.). We will challenge this through a material-informatic analysis, as the basis of this probability is not trivial. See, e.g., [Ramstead et al., 2023], on the introduction of a “formal language for modelling the constraints, forces, potentials, and other quantities determining the dynamics of such systems, especially as they entail dynamics on a space of beliefs (i.e., on a statistical manifold)” for more background on the possible bridges between material and informatics in terms of Bayesian mechanics. Also note that the translation of “probability” in Weber has also been that of “ability”, “capacity”, and other terms, we take one of the most popular formulations here which also aids our purpose of framing this in terms of a probabilistic mechanics. probabilities. We will present novel understandings of concepts such as scripts and attention, by analyzing power in terms of the physics of beliefs [Ramstead et al., 2023], and the metaphysics of social organization, where metaphysics is understood as the possibilistic exploration of state spaces (i.e., the future).
@@
The outline of the article is as follows: in the subsections that follow, we present a necessarily broad and brief elucidation of different understandings of power, primarily by way of philosophy and sociology, in order to set the focus on the concept of power as the control of attention via explorative and exploitative scripts, which can exist as shared metacognitive computing conditions (i.e., agents are relatively aware of the implications of scripts), coercive domination (where agents are subsumed under a narrative which does not benefit them), or: often an inevitable combination of both. In section 2, we explain the formal context of Active Inference as the theoretical ground which will sustain our analysis. In section 3, we synthesize the previous sections towards an explanation of social scripts and power, and how scripts can frame the concept of narratives, attention, coercion and resistance. In section 4, we look at power through the lens of Active Inference, and in section 5, we treat the formal framework integration. We conclude by presenting attentional strategies which can result in alternative modulations of power.
@@ 0.1 Power-pooling in, through and between agents
Gravitational metaphors abound in language [Shepard, 1987, Lakoff and Johnson, 2003, Dennett, 1992, de Jager, 2021, Kent, 2024]333 Lachlan Kent has further argued that the vestibular system, which regulates graviception, also affects the higher-order consciousness of selfhood [Kent, 2024]: gravity constrains physical proprioception and action, as well as the very sense of a narrative self [Dennett, 1992]. ). Such gravitational metaphors find analogues in the network sciences, whereby preferential attachment processes lead to additional power accruing increasingly to those already in possession of it [Barabási and Albert, 1999]. The idea of power being pooled towards complex social attractors was famously presented by Pierre Bordieu, who linked various concepts of capital (cultural, political, etc.) to power as social (at)traction, borrowing the metaphor from physical field theory [^4] Just as electromagnetic fields exert forces on particles depending on their position and properties, Bourdieu conceptualized social fields as spaces where agents’ positions are (pre)determined by their access to different forms of capital (economic, cultural, symbolic, etc.). Bourdieu’s use of field theory therefore frames system structure and the dynamics of agency: a given field structures possible actions by defining what counts as accepted or legitimate, and habitus as inclinations; dispositions, determine how fields are navigated: possibilities. The ensuing social practices herein are the result of heuristics comparable to satisficing/fast and frugal heuristics, which in our context are formalizable through AIF.. We expand on his legacy which takes the ontoepistemological and the empirical to be inseparable—and his desire to ground sociological analyses of habitus in a scientifically tractable manner—by framing these attentional attractors that move systems toward specific configurations, through Active Inference (AIF). Another key thinker of power in the 20th century was Hannah Arendt. Arendt’s relational idea of collective power frames its essence as communicative, public and permanent potential: people may gain momentum together, but power is never guaranteed. Coercion, (symbolic) violence and oppressive authority, in her framing, do not constitute power but are rather reductions of its possibility. In our context, this differentiation would imply the first presentation being shared cultural narratives towards explorative change, and the second as top-down exploitative practices which reduce the possibility of novelty in given structures (systems such as communities; identities, histories, etc.). Similarly, more recently, Grace Blakeley has analyzed the disempowerment of collective political imagination by presenting this as the disillusion555 In our terms, perhaps, understood as the loss of attentional momentum. with shared narratives: democratic engagement in the context of, e.g., British politics, has been affected by a sense of lack of agency (from “there is no alternative” all the way to the not-so “free” market666“This sense of unfreedom is grounded in the deep disparities of power that exist within capitalist societies—many of which are completely invisible. Most people are denied autonomy over their lives, yet we are consistently told that we are free to choose how we live. Life under capitalism means life under a system in which decisions about how we work, how we live, and what we buy have already been made by someone else. Life under capitalism means living in a planned economy, while being told that you are free.” ([Blakeley, 2024], p. 6, our emphasis in bold).). This resulting political disenchantment was framed earlier by Mark Fisher in terms of the privatization or individualization of systemic stress [Fisher, 2009]: through strategic narratives which set public attention on individual responsibility, publics become overwhelmed by having to resolve, at the individual level, stress which is in fact systemic. Whether we are talking about a metastabilizing “invisible hand” or modeling swarm-organization [Wareham et al., 2023]: we know of many systems that have infelicitous critical points when scaling: requisite variety demands that control systems maintain sufficient internal complexity (entropy) to match the complexity of what they aim to control. This can be understood at a number of levels (in modeling: computational tractability problems regarding communication between agents that have to do with temporal tracking and issues such as latency, for example), but at the level of social dynamics and the (in)efficiencies of power asymmetries, this is most obvious in the context of the power of attention: the assertion of a specific narrative, if it gains traction as a collective belief, will result in the simplification (and mobilization) of desire. That is: a strong-enough narrative can become a script which agents incorporate as part of their generative model, rather than as a meta-consideration (a highly possibilistic variable), existing on a dialogical landscape. In crude terms: the difference between the first and the second can be exemplified by the situation where an agent believes that any single solution can solve national scale problems, such as the deportation of immigrants supposedly solving every socio-economic issue in a very large nation. Our framing allows for a new take on the dialectics between the individual and the collective. In terms of questions of information-processing capacity: not only are resource-distribution and the division of labor a systemic fact, but also: since the modulation and projection of future possibilities cannot be computed at the level of the individual due to its high complexity, this is how and why the minimization of uncertainty for social systems is, in fact, social. Agents seek alliances and policies which ensure cognition (i.e., information processing and future-projection) can be distributed [Vasil et al., 2020], as this allows for efficient practical traction because of “epistemic confidence in knowing, interpreting, and acting together” [Shteynberg et al., 2023]. This is the basis for any division of labor, where specialized roles and collaborative synergy optimize task execution. As demonstrated by [Ispolatov et al., 2012], division of labor in biological systems improves fitness by compartmentalizing incompatible processes, while [Labella, 2007] shows its application in robotics, enhancing adaptability and resource efficiency. [Marzband et al., 2017] show that coalition-based strategies in distributed energy resources maximize collective benefits, offering a model for scalable and cooperative systems in human and artificial contexts. This is in contrast with individualized agency—which is a prerequisite for social coercion—where one’s estimation of “divergence in mental states between self and other(s)” allows for the “prediction and interpretation of others’ behavior” [Shteynberg et al., 2023], resulting in shorter-range practical effects. Remaining politically-agnostic in any analysis presents us with some questions: what can we assume to actually know about each other when we act collectively? The concept of a ‘self’ as a single agent seems particularly questionable, especially as all selves exist within a regime of constraints which exert effects on their supposed autonomy. A famous response to the Foucauldian take on power and the individual came from Gilles Deleuze. His concept of the “-dividual” replaces the image of, e.g., the factory worker as a subject of disciplinary capitalism, with the image of a distributed subjectivity, e.g., the owner of a credit card as a subject of libertarian debt, occupying various spatiotemporal scales and modulated not by clock-time or physical enclosure, but by probabilities and contractual fixation/limitation [Deleuze, 1992]. The latter condition emerges, Deleuze shows, as a result of novel paradigms of technocapitalism which present new constraint regimes. Distribution, collectivity, or shared sameness, therefore, do not always equal autonomous collective traction, but can also be the result of policies which pool power towards attractors, where the agents implicated within them may not even be aware of their implication (the examples in capitalism are vast: we have very little idea of the complexity behind the products we buy into, but we are coerced into buying nonetheless [^7] For more on this theme, see “Post-control Script-societies” [de Jager, 2023].). In the interest of this article, the concept tying all these processes towards an analysis of power dynamics is that of attention. We will briefly treat this below (and in later sections) before moving to section 2 for an exposition of the AIF framework.
@@ 0.2 Attention as power: shared narratives and awareness or coercion
Exerting social control over others’ actions begins at the level of attention: whether this be by attracting attention (“follow me!”), or directing attention (“look at that!”). Our interest in formalizing certain aspects of power dynamics is directing attention towards (metacognitive and collective) scripts with the capacity of empowering responses to counter dominant strategies which, currently, minimize uncertainty for the very few [^8] Again, often by allocating too much importance to the atomized individual, which is a predictably convenient way to simplify the actions of complex distributed processes such as individuals, into overseeable patterns. There is nothing essentially “wrong” with this, this is the way perception tracks reality. However, revealing these dynamics allows us an entrance into the ways in which we may modulate (the effects of) these phenomena/concepts.. Shared structures of meaning can exist as collective metacognitive computing conditions, i.e., conditions in which agents are—relatively—aware of the implications of the scripts they follow, or coercive domination: where agents are subsumed under a script which, often, does not benefit their uncertainty-minimization on the long run. If we understand successful narratives as those which exert the most effect [^9] Often “fast and frugal” [Gigerenzer, 2008, Simon, 1978] heuristics and narratives, which are low-risk and, in most political contexts, imply promises or solutions impossible to accomplish: “you have been lied to”, “the other is the problem”, etc., analyzing them in terms of attention seems logical: in the context of uncertainty-minimization, simple narratives have the capacity to dominate vast regimes of attention (economies), as they are scripts which minimize the complexities of life into low-computation, predictable visions [^10] Whether these are dialogical and representational/symbolic or surreptitious and encrypted: we’re dealing with different strategies, but similar results..
@@
In the transition between an economy centered around material and spatial capital, towards one of temporal, speculative capital, we can straightforwardly connect attention to computation by understanding the implicit link between risk-assessment and investment: (possible) future gains ought to be calculated, and directing investment begins with attentional gains111111“A sufficiently large financial institution has the power to direct investment into certain technologies and therefore to determine which futures are available and which ones are foreclosed. And all these firms can consolidate their power by buying up competitors, erecting barriers to entry, and crushing workers’ attempts to organize, further insulating them from competitive pressure and giving them significant authority both within their domains and over society as a whole.” [Blakeley, 2024].. Beyond the strictly financial realm, in the context of futurology and planetary survival, the concept of AI technologies as “powerful” grounds the arguments surrounding possible existential risks. E.g., Cappelen et al. rely on the concept of “power” in order to assess to what extent humans may or may not lose control over their own technological creations [Cappelen et al., 2024]. Likewise, other authors make claims for “powerful yet narrow AI systems” as incapable of accommodating the semantic complexity which grounds mortal life [Ororbia and Friston, 2023] [^12] They continue: “One might view the future as residing in not viewing intelligence as a difference engine but instead as an artificial form of sentience: one that is capable of self-healing and self-repairing with autonomy — its existential imperative being to persist in (generalized) synchrony with its world.” (ibid.). The questions of “autonomy” and “synchrony” are precisely the ones we are after in this article, by observing how power, in mortally-determined social systems, is driven by attention.. What these assessments have in common is a concern with attention as a resource: a collectively predicted future is one where attention is directed towards synchronized goals (whether this be investment, alignment, etc.).
@@
Ultimately, because potential power (as attention) is both a material and an informational force, “seizing the means of computation” (Doctorow, 2023) follows in line with a Marxist analysis of capital in our moving paradigm: from materialist analyses of territorial and labor time domination, through a semantic-sociocultural analysis [Bourdieu, 2018], to informational analyses where we ought to think about the production of the future by means of projective computation. Power is thus based on the controlling of attention by way of scripts (again: narratives, representation, identities, symbolisms, etc.) and its current substrate in our contemporary landscape is computation: both in the flesh and in silico. As we will see in section 3, power’s dual nature (as force attraction and information computation) in the social realm thus implies the use of scripts as they are the means by which agents gain or lose the ability to direct the attention of other agents (with access to desired policies), and the ability to effectively process information in favor of these policies (and other longer-term possibilities).
@@
The variables modulating regimes of attention that have practical modeling potential are plenty, and in sections 4 and 5, our interest is in the control, influence, and synchronization of attention, which lead to social power structures. Our focus will be on empowerment quantified as agents’ potential influence over their environment (which is tracked by assessing the mutual information between actions and states). We will analyze this through formal methods that are tractable, but this presents us with indications of future modeling possibilities, which in our context remain unformalized as they are practically intractable. We can think, for example, about the modeling of “inheritance” or “provenance” variables: wherever/however an agent exists, they could be born into a rich family, in a specific part of the world, etc., and these factors have an influence on how others interpret them (and therefore on life-expectancy: mortality, and many other crucial effects). One can also posit a possible “frictions/contingencies/contextualities” variable—in colloquial terms “you have to work with what you’ve got”—by which, even if agents can attract others’ attention and process information effectively towards the effectuation of a specific policy, the projected goal and thus possible state space is constrained by a context which can be highly divergent from the “ideal” one they project as required for bringing a goal about. Confidence in policies is a well-framed and studied variable in the context of AIF, and in the literature it is often directly linked to the concept of power: “self-confidence [is] the person’s experience of their power to act in the world” [Kiverstein et al., 2019]. There are more, of course, and it depends on how granular one may want to go, these are examples which give the reader an indication of our intentions in later sections.